Conversion of decimal to degree or degree to decimal.
Problem.
50° 1/2' 1/4"
Thanks.

Answer by lwsshak3(11628)

Conversion of decimal to degree or degree to decimal.
Problem.
convert degree to decimal:
50° 1/2' 1/4" =50˚+30"+1/4"=50˚+30.25"=50+(30.25/3600)≈50.0084˚
note:
1deg=60 min
1min=60 sec
1deg=3600 sec